OpenStudy (anonymous):

If the sum of the legs of a right triangle (not including the hypotenuse) is 49 inches and the hypotenuse is 41 inches, find the area of the triangle.

OpenStudy (amoodarya):

name them as a,b,c a= hypotenuse so b+c=49 a=41 and \[a^2=b^2+c^2\] now you have to solve the system of equation

OpenStudy (anonymous):

You need to find out what x and y are and then multiply them together, and then divide the answer by 2. This is because the area of a right triangle is (1/2)(base*height). The hypotenuse is 41. That is h. The sum of the legs is x + y = 49. x^2 + y^2 = h^2

OpenStudy (amoodarya):

b+c=49 so find (b+c)^2=b^2+c^2 +2bc so 2bc +a^2=(c+b)^2 2bc +41^2=49^2 and you need to find area=1/2 bc

OpenStudy (anonymous):

Ohhh. I was squaring the wrong thing.